A frictionless wire AB is fixed on a sphere of radius R. A very small spherical ball slips on this wire. The time taken by this ball to slip from A to B is?

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
time taken by this ball to slip from A to B is:
AB = R/(cos theta) theta = 45o                            
AB = √2 R = 2R/√2 = 2R cos?
  • 2R cos? = 1/2x gcos? x t2  
t = 2 √R/g
Regards
Arun (askIITians forum expert)